# MEAL: A Benchmark for Continual Multi-Agent Reinforcement Learning

### v1

@@ 1 Introduction
Continual RL has recently attracted growing interest [13, 7, 8, 11], but remains largely unexplored in multi-agent settings [34, 35]. Combining the two introduces unique challenges. In cooperative environments, agents must establish implicit conventions or roles for effective coordination [28]. As tasks or dynamics shift, these conventions can break down, making continual MARL significantly harder than its single-agent counterpart. Forgetting past partners or roles can cause the entire team to fail, amplifying the impact of catastrophic forgetting through inter-agent dependencies. Unlike traditional MARL, CMARL involves non-stationarity not only due to the presence of other learning agents, but also from a shifting task distribution [35]. This dual pressure demands agents that can generalize, adapt, and transfer knowledge more robustly than in standard single-agent continual or static multi-agent settings. This setting is relevant for applications where agents must adapt to evolving environments without forgetting prior coordination strategies. For instance, autonomous vehicles must navigate unseen roads, adapt to new traffic regulations, and interact with unfamiliar human drivers, while occasionally coordinating with other AVs. Similarly, warehouse robots deployed in a new facility must quickly adapt to different layouts and workflows, while preserving established collaborative behaviors.
To analyze how current methods handle the interplay between CL and MARL, and to drive progress in this domain, we introduce MEAL, the first benchmark for CMARL. To the best of our knowledge, MEAL [^1] The code and environments are accessible on GitHub. is also the first continual RL library to leverage JAX for end-to-end GPU acceleration. Traditional CPU-based benchmarks are limited to short sequences (5–15 tasks) due to low environment throughput and task diversity [27, 23, 30], making them ill-suited for the computational demands of CL across long task sequences. MEAL’s end-to-end JAX pipeline removes this barrier, enabling training on up to 100 tasks within a few hours on a single desktop GPU. This unlocks new research directions for scalable, cooperative continual learning in resource-constrained settings.
MEAL is built on Overcooked [6], a widely used cooperative MARL environment [15, 32, 28] that provides a strong foundation for benchmarking. Prior work has shown that agents tend to exploit spurious correlations in fixed layouts, resulting in poor generalization even under minor modifications [17]. This makes Overcooked particularly well-suited for continual learning: even small layout variations can present a significant challenge. To succeed across a sequence of such tasks, agents must avoid overfitting to layout-specific behaviors and instead learn coordination strategies that are robust and transferable.
@@
CORA [23] 31 Mixed
MPE [21] 7 Continuous
SMAC [26] 14 Discrete
Continual World [31] 10 Continuous
Melting Pot [2] 49 Discrete
Google Football [18] 14 Discrete
JaxMARL [25] 33 Mixed
COOM [30] 8 Discrete
@@
The contributions of our work are three-fold. (1) We introduce MEAL, the first CMARL benchmark, consisting of handcrafted and procedurally generated Overcooked environments spanning three difficulty levels. (2) We leverage JAX to build the first end-to-end GPU-accelerated task sequences for continual RL, enabling efficient training on low-budget setups. (3) We implement five popular CL methods in JAX and evaluate them in various MEALs, revealing key shortcomings in retaining cooperative behaviors and adapting to shifting roles across tasks.
@@ 2 Related Work
Continual reinforcement learning studies how agents can learn sequentially from a stream of tasks without forgetting previous knowledge. A wide range of methods have been proposed, including regularization-based approaches such as EWC [16], SI [36], and MAS [3]; architectural strategies such as PackNet [20]; and replay-based methods like RePR [4]. More recent works focus on scalability [13], memory efficiency [8], and stability during training [7]. However, these methods are almost exclusively developed for single-agent settings, and their behavior under multi-agent coordination remains largely unexplored.
@@
In MARL, multiple agents learn to act in a shared environment, often with partial observability and either cooperative or competitive goals [14, 22]. A major focus has been on cooperative settings, where agents share a reward function and must learn to coordinate [19, 12]. Popular algorithms include IPPO [9], VDN [29], QMIX [24], and MAPPO [33]. Many benchmarks assume a static environment and fixed task, making them unsuitable for studying continual learning or transfer across environments.
@@
Standard CRL benchmarks include Continual World [31], COOM [30], and CORA [23]. While effective in single-agent settings, they either lack multi-agent capabilities or suffer from slow CPU-bound environments. For MARL, environments like SMAC [26], MPE [21], and Melting Pot [2] are widely used, but are not designed for continual evaluation. Overcooked [6] has emerged as a useful domain for studying coordination, with recent implementations in JAX [25]. Our benchmark builds on Overcooked and introduces procedural variation to create long task sequences for continual MARL.
@@
The Overcooked environment [6] is a cooperative multi-agent benchmark inspired by the popular video game of the same name. Agents control chefs in a grid-based kitchen, coordinating to prepare and deliver dishes through sequences of interactions with environment objects such as pots, ingredient dispensers, plate stations, and delivery counters. The environment is designed to require both motion and strategy coordination, making it a standard testbed for evaluating collaborative behaviors. Compared to the large state spaces and high agent counts in benchmarks like Melting Pot [2] and SMAC [26], Overcooked operates on small grid-based environments with only two agents. However, its complexity arises not from scale but from credit assignment challenges due to shared rewards, and the need for precise coordination, as agents must execute tightly coupled action sequences to complete tasks successfully [14]. However, its fully observable and symmetric setup reduces the need for explicit communication.
@@ 3 Preliminaries
We formulate the setting as a fully observable cooperative multi-agent task, modeled as a Markov game defined by the tuple $\langle N,S,{A^{i}}_{i\in N},P,R,\gamma\rangle$, where $N$ is the number of agents, $S$ is the state space, $A^{i}$ is the action space of agent $i$ with joint action space $A=A^{1}\times\dots\times A^{N}$, $P:S\times A\times S\rightarrow[0,1]$ is the transition function, $R:S\times A\times S\rightarrow\mathbb{R}$ is a shared reward function, and $\gamma\in[0,1)$ is the discount factor. In the fully observable setting, each agent receives the full state $s\in S$ at every time step.
@@
We consider a continual MARL setting in which a shared policy $\pi_{\theta}={\pi^{i}_{\theta}}_{i\in N}$ is learned over a sequence of tasks $\mathcal{T}={\mathcal{M}_{1},\dots,\mathcal{M}_{T}}$, where each $\mathcal{M}_{t}=\langle N,S_{t},{A^{i}}{i\in N},P_{t},R_{t},\gamma\rangle$ is a fully observable cooperative Markov game with consistent action and observation spaces. At training phase $t$, agents interact exclusively with $\mathcal{M}_{t}$ for a fixed number of iterations $\Delta$, collecting trajectories ${\tau_{t,1},\dots,\tau_{t,\Delta}}$ to update their policy. Past tasks and data are inaccessible, and no joint training or replay is allowed. The focus of this work is on the task-incremental setting, where the task identity is known during training but hidden at evaluation. The objective is to maximize cumulative performance across all tasks and mitigate forgetting.
@@ 4 MEAL
We introduce MEAL, the first benchmark for CMARL, built on the JaxMARL [25] version of Overcooked. JAX [5] provides just-in-time compilation, automatic differentiation, and vectorization through XLA, enabling high-performance and accelerator-agnostic computation. We incorporate the original five layouts from Overcooked-AI [6] and design 20 additional handcrafted environments.
4.1 Environment Dynamics
@@
Agents receive a shared team reward: $r_{t}=r_{\text{deliver}}+r_{\text{onion}}\cdot\mathbbm{1}_{\{\text{onion\_in\_pot}\}}+r_{\text{plate}}\cdot\mathbbm{1}_{\{\text{plate\_pickup}\}}+r_{\text{soup}}\cdot\mathbbm{1}_{\{\text{soup\_pickup}\}}$, where $r_{\text{deliver}}=20$ is the reward for delivering a cooked soup, and the remaining terms provide shaped rewards for intermediate progress. We include two reward settings: in the sparse setting, $r_{\text{onion}}=r_{\text{plate}}=r_{\text{soup}}=0$; in the dense setting, $r_{\text{onion}}=r_{\text{plate}}=3$, and $r_{\text{soup}}=5$.
Score Function
We measure performance as the total number of soups successfully delivered by the team during an episode. Since MEAL layouts vary in size, structure, and the number of interactive elements, the maximum number of soups that can be delivered in a fixed time horizon differs across tasks. To allow fair comparisons across sequences and environments, we normalize returns using a programmatically computed performance estimate. For each layout we estimate the fastest possible single-agent cook–deliver cycle, accounting for the shortest paths between onion piles, pots, plate piles, and delivery counters, the fixed cooking time, and pickup/drop interactions. Repeating this cycle over the episode length yields a soup total estimate. This ratio equals 1 when an agent matches the single-agent optimum and can exceed 1 if multiple agents collaborate more efficiently than our single-agent model assumes. A detailed description can be found in Appendix A.1.
@@ 4.2 MEAL Generator
(a) Level 1 (Easy): $6\leq$ width/height $\leq 7$, obstacle density $\approx 15\%$. Layouts are compact, making exploration easy. Interactable items are close together, and travel distances are short. Agents can often complete the task independently with minimal coordination.
(b) Level 2 (Medium): $8\leq$ width/height $\leq 9$, obstacle density $\approx 25\%$. Exploration becomes harder as key items are more spread out. Layouts often introduce chokepoints, requiring agents to coordinate movement and avoid congestion.
(c) Level 3 (Hard): $10\leq$ width/height $\leq 11$, obstacle density $\approx 35\%$. Layouts are more likely to split the map into disjoint regions, forcing agents to specialize. Solving the task requires deliberate cooperation and division of labor.
@@
Existing continual RL benchmarks only provide a fixed set of tasks [27, 23, 30]. To avoid over-fitting to a fixed set of environments, we procedurally generate new Overcooked kitchens on the fly. The generator draws a random width and height from the specified range, places an outer wall, then sequentially injects the interactive tiles (goal, pot, onion pile, plate pile), extra internal walls to match the target obstacle density, and finally, the agents’ starting positions. Figure 2 depicts the pipeline, and the process is described more in-depth in Appendix A.2. Each candidate grid is accepted only if a built-in validator module confirms that both agents can complete at least one cook–deliver cycle. This yields a continuous space of solvable, variable-sized kitchens that we can learn continually. We bring further details about the validator in Appendix A.3. Our approach offers a virtually infinite supply of tasks and evaluates true lifelong learning under continual exposure to unseen configurations. To ensure reproducibility and a fair comparison between methods, the generation process can be fully controlled via a user-specified random seed.
4.3 Layout Difficulty
We categorize environment difficulty based on procedurally generated layout characteristics. We vary the (1) grid width, (2) grid height, and (3) obstacle density. This approach produces diverse spatial configurations while maintaining consistent difficulty within each level. Figure 1 illustrates representative layouts for each difficulty tier. As grid size and the number of impassable tiles increase, agents must develop more sophisticated coordination strategies. Higher difficulty layouts feature longer paths between key items, tighter bottlenecks, and greater structural variability, all of which make exploration, retention, and adaptation more challenging. Appendix C.1 presents a performance comparison across difficulty levels. Level 1 tasks are designed such that existing methods can achieve reasonably high scores, enabling better comparisons and behavioral analysis. Higher levels are intended to challenge future methods. Although we currently include three difficulty levels, it is straightforward to extend the framework. While obstacle density has a practical upper bound, grid size can be increased arbitrarily to scale up environment complexity.
(a) Empty grid with outer walls
(b) Interactive stations randomly placed
(c) Grid filled with walls to match obstacle density
(d) Agents added & unreachable tiles pruned
@@ 4.4 Task Sequences
Rather than a continuous domain shift, MEAL sequences involve discrete task boundaries, where agents transition between clearly distinct environments. By default, the task ID is accessible, aligning with the task-incremental learning paradigm.
@@ 4.5 Evaluation Metrics
We evaluate methods on three core metrics: Average Performance, Forgetting, and Plasticity. Let $s_{i}(j)$ denote the normalized score (see Section 4.1) on task $j$ after training on task $i$, and let the task sequence consist of $N$ tasks.
Average Performance
We define average performance as the mean normalized score across all tasks at the end of training. This metric captures the balance between forward transfer and retention:
$\begin{array}[]{c}\mathcal{A}=\frac{1}{N}\sum_{i=1}^{N}s_{N}(i)\end{array}\vspace{-0.75em}$ (1)
@@
Forgetting quantifies the degradation in performance on past tasks due to interference from training on later ones. For each task $i<N$, we compute the difference between the performance immediately after training and at the end of the sequence:
$\begin{array}[]{c}\mathcal{F}=\frac{1}{N-1}\sum_{i=1}^{N-1}\left(s_{i}(i)-s_{N}(i)\right)\end{array}\vspace{-0.75em}$ (2)
Plasticity
To evaluate continual training capacity over long task sequences, we measure the model’s ability to fit new tasks under continual learning constraints. We compute the average training score (i.e., final performance on the current task right after training) across the entire sequence:
$\begin{array}[]{c}\mathcal{P}=\frac{1}{N}\sum_{i=1}^{N}s_{i}(i)\end{array}\vspace{-0.75em}$ (3)
This isolates how quickly and effectively the method learns a new task under capacity constraints, independently of retention. Unlike Average Performance, Plasticity does not require evaluation on previously seen tasks.

### v2

@@ 1 Introduction
Continual RL has recently attracted growing interest Hafez and Erekmen (2024); Chen et al. (2024); Chung et al. (2024); Erden et al. , but remains largely unexplored in multi-agent settings Yuan et al. (2023, 2024). Combining the two introduces unique challenges. In cooperative environments, agents must establish implicit conventions or roles for effective coordination Strouse et al. (2021). As tasks or dynamics shift, these conventions can break down, making continual MARL significantly harder than its single-agent counterpart. Forgetting past partners or roles can cause the entire team to fail, amplifying the impact of catastrophic forgetting through inter-agent dependencies. Unlike traditional MARL, CMARL involves non-stationarity not only due to the presence of other learning agents, but also from an unknown task distribution Yuan et al. (2024). This dual pressure demands agents to generalize, adapt, and transfer knowledge more robustly than in standard single-agent continual or static multi-agent settings. This setting is prevalent in applications where agents must adapt to evolving environments without forgetting prior coordination strategies. For instance, autonomous vehicles must navigate unseen roads, adapt to new traffic regulations, and interact with unfamiliar human drivers, while coordinating with other AVs. Similarly, warehouse robots deployed in a new facility must quickly adapt to unseen layouts and workflows, while preserving established collaborative behaviors.
To analyze how current methods handle the interplay between CL and MARL, and to drive progress in this domain, we introduce MEAL, the first benchmark for CMARL. To the best of our knowledge, MEAL [^1] The code and environments are accessible on GitHub. is also the first continual RL library to leverage JAX for end-to-end GPU acceleration. Traditional CPU-based benchmarks are limited to short sequences (5–15 tasks) due to low environment throughput and task diversity Sorokin and Burtsev (2019); Powers et al. (2022); Tomilin et al. (2023), making them ill-suited for the computational demands of CL across long task sequences. MEAL’s end-to-end JAX pipeline removes this barrier, enabling training on up to 100 tasks within a few hours on a single desktop GPU. This unlocks new research directions for scalable, cooperative continual learning in resource-constrained settings.
MEAL is built on Overcooked Carroll et al. (2019), a widely used cooperative MARL environment Hu et al. (2020); Wu et al. (2021); Strouse et al. (2021), providing a strong foundation for benchmarking. Prior work has shown that agents tend to exploit spurious correlations in fixed layouts, resulting in poor generalization even under minor modifications Knott et al. (2021). This makes Overcooked particularly well-suited for continual learning: even small layout variations can present a significant challenge. To succeed across a sequence of such tasks, agents must avoid overfitting to layout-specific behaviors and instead learn coordination strategies that are robust and transferable.
@@
CORA Powers et al. (2022) 31 Mixed
MPE Mordatch and Abbeel (2018) 7 Continuous
SMAC Samvelyan et al. (2019) 14 Discrete
Continual World Wołczyk et al. (2021) 10 Continuous
Melting Pot Agapiou et al. (2022) 49 Discrete
Google Football Kurach et al. (2020) 14 Discrete
JaxMARL Rutherford et al. (2024) 33 Mixed
COOM Tomilin et al. (2023) 8 Discrete
@@
The contributions of our work are three-fold. (1) We introduce MEAL, the first CMARL benchmark, consisting of procedurally generated Overcooked environments spanning three difficulty levels. (2) We leverage JAX to build the first end-to-end GPU-accelerated task sequences for continual RL, enabling efficient training on low-budget setups. (3) We implement six popular CL methods in JAX and evaluate them in MEAL, revealing key shortcomings in retaining cooperative behaviors and adapting to shifting roles across tasks.
@@ 2 Related Work
CRL studies how agents can learn sequentially from a stream of tasks without forgetting previous knowledge. A wide range of methods have been adapted from the CL literature to facilitate the RL setting, including regularization-based approaches such as EWC Kirkpatrick et al. (2017), SI Zenke et al. (2017), and MAS Aljundi et al. (2018); architectural strategies such as PackNet Mallya and Lazebnik (2018); and replay-based methods like RePR Atkinson et al. (2021). More recent works focus on scalability Hafez and Erekmen (2024), memory efficiency Chung et al. (2024), and stability during training Chen et al. (2024). However, these methods are almost exclusively developed for single-agent settings, and their behavior under multi-agent coordination remains largely unexplored.
@@
In MARL, multiple agents learn to act in a shared environment, often with partial observability and either cooperative or competitive goals Hernandez-Leal et al. (2019); OroojlooyJadid and Hajinezhad (2019). A major focus has been on cooperative settings, where agents share a reward function and must learn to coordinate Lowe et al. (2017); Foerster et al. (2018). Popular algorithms include IPPO De Witt et al. (2020), VDN Sunehag et al. (2017), QMIX Rashid et al. (2020), and MAPPO Yu et al. (2022). Many benchmarks assume a static environment and fixed task, making them unsuitable for studying continual learning or transfer across environments.
@@
Standard CRL benchmarks include Continual World Wołczyk et al. (2021), COOM Tomilin et al. (2023), and CORA Powers et al. (2022). While effective in single-agent settings, they either lack multi-agent capabilities or suffer from slow CPU-bound environments. For MARL, environments like SMAC Samvelyan et al. (2019), MPE Mordatch and Abbeel (2018), and Melting Pot Agapiou et al. (2022) are widely used, but are not designed for continual evaluation. Overcooked Carroll et al. (2019) has emerged as a useful domain for studying coordination, with recent implementations in JAX Rutherford et al. (2024). Our benchmark builds on Overcooked and introduces procedural variation to create long task sequences for continual MARL.
@@
The Overcooked environment Carroll et al. (2019) is a cooperative multi-agent benchmark inspired by the popular video game of the same name. Agents control chefs in a grid-based kitchen, coordinating to prepare and deliver dishes through sequences of interactions with environment objects such as pots, ingredient dispensers, plate stations, and delivery counters. The environment is designed to require both motion and strategy coordination, making it a standard testbed for evaluating collaborative behaviors. Compared to the large state spaces and high agent counts in benchmarks like Melting Pot Agapiou et al. (2022) and SMAC Samvelyan et al. (2019), Overcooked operates on small grid-based environments with only two agents. However, its complexity arises not from scale but from credit assignment challenges due to shared rewards, and the need for precise coordination, as agents must execute tightly coupled action sequences to complete tasks successfully (Hernandez-Leal et al., 2019).
@@ 3 Preliminaries
We formulate the setting as a fully observable cooperative multi-agent task, modeled as a Markov game defined by the tuple $\langle N,S,A,P,R,\gamma\rangle$, where $N$ is the number of agents, $S$ is the state space, $A^{i}$ is the action space of agent $i$ with joint action space $A=A^{1}\times\dots\times A^{N}$, $P:S\times A\times S\rightarrow[0,1]$ is the transition function, $R:S\times A\times S\rightarrow\mathbb{R}$ is a shared reward function, and $\gamma\in[0,1)$ is the discount factor. In the fully observable setting, each agent receives the full state $s\in S$ at every time step.
@@
We consider a continual MARL setting in which a shared policy $\pi_{\theta}={\pi^{i}_{\theta}}_{i\in N}$ is learned over a sequence of tasks $\mathcal{T}={\mathcal{M}_{1},\dots,\mathcal{M}_{N}}$, where each $\mathcal{M}_{i}=\langle N,S_{i},A,P_{i},R_{i},\gamma\rangle$ is a fully observable cooperative Markov game with consistent action and observation spaces. At training phase $i$, agents interact exclusively with $\mathcal{M}_{i}$ for a fixed number of iterations $\Delta$, collecting trajectories $\tau_{i,1},\dots,\tau_{i,\Delta}$} to update their policy. Past tasks are inaccessible, and no joint training is allowed. The objective is to maximize performance on all tasks in the sequence.
@@ 4 MEAL
We introduce MEAL, the first benchmark for CMARL, built on the JaxMARL Rutherford et al. (2024) version of Overcooked. JAX Bradbury et al. (2018) provides just-in-time compilation, automatic differentiation, and vectorization through XLA, enabling high-performance and accelerator-agnostic computation.
4.1 Environment Specifications
Dynamics
Agents act synchronously at each time step. Moves into walls or occupied tiles are no-ops, and simultaneous swaps are disallowed (both agents remain in place). Agents can interact with the tile they are facing, which deterministically updates the object’s state (pick/place, add onion, plate, deliver). Pots initiate a fixed cook timer of $c_{\text{cook}}{=}20$ steps when the third onion is added, and the cooked soup can only be plated upon completion.
@@ 4.1 Environment Specifications
Agents receive a shared team reward: $r_{t}=r_{\text{deliver}}+r_{\text{onion}}\cdot\mathbbm{1}_{\{\text{onion\_in\_pot}\}}+r_{\text{plate}}\cdot\mathbbm{1}_{\{\text{plate\_pickup}\}}+r_{\text{soup}}\cdot\mathbbm{1}_{\{\text{soup\_pickup}\}}$, where $r_{\text{deliver}}=20$ is the reward for delivering soup, and the other terms provide shaped rewards for intermediate progress. We include two reward settings: in the sparse setting, $r_{\text{onion}}=r_{\text{plate}}=r_{\text{soup}}=0$; in the dense setting, $r_{\text{onion}}=r_{\text{plate}}=3$, and $r_{\text{soup}}=5$.
@@ 4.2 MEAL Generator
(a) Empty grid drawn with outer walls.
(b) Interactive stations sampled at random locations.
(c) Grid filled with walls to match obstacle density.
(d) Agents added and unreachable tiles pruned.
Existing continual RL benchmarks only provide a fixed set of tasks Sorokin and Burtsev (2019); Powers et al. (2022); Tomilin et al. (2023). To avoid over-fitting to a fixed set of environments, we procedurally generate new Overcooked kitchens on the fly. The generator $G$ draws a random width and height from the specified range, places an outer wall, then sequentially injects the interactive tiles (goal, pot, onion pile, plate pile), extra internal walls to match the target obstacle density, and finally, the agents’ starting positions. Figure 1 depicts the steps in the pipeline, and the process is described more in-depth in Appendix A.2. Each candidate grid is accepted only if a built-in validation module confirms that both agents can complete at least one cook–deliver cycle. This yields a continuous space of solvable, variable-sized kitchens that we can learn continually. We bring further details about the validator in Appendix A.3. Our approach offers a virtually infinite supply of tasks and evaluates true lifelong learning under continual exposure to unseen configurations. To ensure reproducibility and a fair comparison between methods, the generation process can be fully controlled via a user-specified random seed.
4.3 Layout Difficulty
We categorize environment difficulty based on procedurally generated layout characteristics. We vary the (1) grid width, (2) grid height, and (3) obstacle density. This approach produces diverse spatial configurations while maintaining consistent difficulty within each level. Figure 2 depicts layouts of each difficulty. As grid size and the number of impassable tiles increase, agents must develop more sophisticated coordination strategies. Higher difficulty layouts feature longer paths between key items, tighter bottlenecks, and greater structural variability, all of which make exploration, retention, and adaptation more challenging. Level 1 tasks are designed to be simple enough for existing methods to achieve reasonably high scores, enabling better comparisons and behavioral analysis. Higher levels are intended to challenge future methods. Although we currently include three difficulty levels, it is straightforward to extend the framework. Although obstacle density has a practical upper bound, the grid size can be increased arbitrarily to scale up environment complexity.
(a) Level 1 (Easy): $6\leq$ width/height $\leq 7$, obstacle density $\approx 15\%$. Layouts are compact, making exploration easy. Interactable items are close together, making travel distances short. Agents can often complete the task independently with no coordination.
(b) Level 2 (Medium): $8\leq$ width/height $\leq 9$, obstacle density $\approx 25\%$. Exploration is harder as stations are more spread out. Layouts often introduce chokepoints, requiring agents to coordinate movement and avoid congestion.
(c) Level 3 (Hard): $10\leq$ width/height $\leq 11$, obstacle density $\approx 35\%$. Layouts are likely to split the map into disjoint regions, forcing agents to specialize. Solving the task requires deliberate cooperation and division of labor.
@@ 4.4 Task Sequences
MEAL provides discrete task sequences $\mathcal{T}=(\mathcal{M}_{1},\dots,\mathcal{M}_{N})$ rather than a continuous domain shift. For a chosen difficulty level $\ell\!\in\!\{1,2,3\}$, we sample $N$ solvable layouts i.i.d. from the generator $G_{\ell}$ with a fixed seed. At task boundaries, we carry over the optimizer state and policy parameters, reset rollout buffers, and advance the RNG. We explore three sequence regimes: (i) fixed-level, where all tasks are drawn from the same difficulty level; (ii) curriculum, where sequences contain an equal number of tasks in increasing difficulty level (see Appendix E), and (iii) repetition, where a fixed sequence is repeated $r$ times (used in the network plasticity study in Section 5.5).
@@ 4.5 Evaluation Metrics
We measure task performance as the number of soups successfully delivered during an episode. Since MEAL layouts vary greatly in size, structure, number of interactive stations, and distances between them, raw delivery counts are not directly comparable. To unify performance across tasks and sequences for continual learning metrics, we divide the delivery count by an upper bound. We calculate this bound for each layout by finding the optimal cook-deliver cycle for a single agent (see Appendix A.1), accounting for the shortest paths between onion piles, pots, plate piles, and delivery counters, the fixed cooking time, and pickup/drop interactions. Repeating this cycle over the length of the episode yields a soup-total count. A normalized score of 1 indicates that the agent(s) achieved the optimal performance possible for a single agent, while values above 1 reflect effective cooperation that exceeds solo efficiency. Let $s_{i}(t)$ denote this normalized delivery score on task $i$ at timestep $t$. Suppose that the training sequence consists of $N$ tasks, each lasting $\Delta$ steps, resulting in a total of $T=N\cdot\Delta$ timesteps. The $i$-th task is therefore trained during the interval $t\in[(i-1)\Delta,i\Delta]$. Following prior work Wołczyk et al. (2021); Tomilin et al. (2023), MEAL measures performance in three continual learning metrics:
Average Normalized Score
To capture the balance between stability and plasticity, we report the mean score across all tasks at the end of training:
$\mathcal{A}=\frac{1}{N}\sum_{i=1}^{N}s_{i}(T)\vskip-7.5pt$ (1)
@@
Forgetting quantifies the degradation in performance on past tasks due to interference from training on later ones. We exclude the final task, as no forgetting can occur without further training. For each task $i<N$, we compute the difference between the average score over the final $k$ steps after training on task $i$ and the average over the final $k$ steps of the full sequence $T$:
$\mathcal{F}=\frac{1}{N-1}\sum_{i=1}^{N-1}\left(\frac{1}{k}\sum_{t=i\Delta-k}^{i\Delta-1}s_{i}(t)-\frac{1}{k}\sum_{t=T-k}^{T-1}s_{i}(t)\right)\vskip-7.5pt$ (2)
Forward Transfer
Forward transfer measures how prior experience accelerates learning of new tasks. Rather than evaluating final performance, it captures how quickly each task is learned relative to a single-task baseline. We first compute the normalized area under the learning curves (AUC) for both the agent and the baseline:
$\mathrm{AUC}_{i}=\frac{1}{\Delta}\int_{(i-1)\Delta}^{i\Delta}s_{i}(t)\,dt,\qquad\mathrm{AUC}^{\mathrm{b}}_{i}=\frac{1}{\Delta}\int_{0}^{\Delta}s^{\mathrm{b}}_{i}(t)\,dt.\vskip-7.5pt$ (3)
Forward transfer for task $i$ is the area between these curves, positive when prior training helps, and negative when it hinders. We report the average forward transfer across the whole sequence:
$\mathcal{FT}_{i}=\frac{\mathrm{AUC}_{i}-\mathrm{AUC}^{\mathrm{b}}_{i}}{1-\mathrm{AUC}^{\mathrm{b}}_{i}},\qquad\mathcal{FT}=\frac{1}{N}\sum_{i=1}^{N}\mathrm{FT}_{i}.\vskip-7.5pt$ (4)